OpenStudy (nuccioreggie):

Malia solved an equation, as shown below: Step 1: 6x = 48 Step 2: x = 48 – 6 Step 3: x = 42 Part A: Is Malia's solution correct or incorrect? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation. (6 points) Part B: How many solutions will this equation have? (4 points)

OpenStudy (whovianchick):

Nevermind to that message I just sent you XD

OpenStudy (nuccioreggie):

its ok lol

OpenStudy (whovianchick):

Malia is incorrect. She subtracted 6, instead of dividing 6. The equation is multiplying 6 by x, so we have to get rid of the 6 by dividing. The correct steps would be: 6x=48 6x/6=48/6 x=8 Also, there's only one solution, which is 8. It's a linear equation, so it wouldn't have 2, and I've proven that it doesn't have 0 solutions. :) Still with me?